# PAFT: Prompt-Agnostic Fine-Tuning

### v1

@@ 5 Empirical Results
In this section, we conduct extensive experiments to evaluate the effectiveness and efficiency of our proposed PAFT framework. We begin by detailing the datasets and experimental setup in Section 5.1, followed by a comprehensive analysis of the main results in Section 5.2. Additionally, we perform ablation studies to investigate the impact of key components of our framework, as discussed in Section 5.3.
@@ 5.1 Datasets and Setup
To evaluate the performance of our proposed PAFT method, we focus on reasoning and reading comprehension tasks, as these domains are particularly susceptible to prompt variations. As PAFT is the first work to address the prompt robustness problem in large language models (LLMs) through training, we generate task-specific candidate prompts for each downstream task. Following the dataset selection process of Hu et al. (2023a); Wei et al. (2024), we select the Winogrande Sakaguchi et al. (2019), PIQA Bisk et al. (2019), and Hellaswag Zellers et al. (2019) reasoning benchmarks and additionally include the RACE Lai et al. (2017) reading comprehension benchmark. These datasets are widely recognized for their ability to assess reasoning and comprehension, provide independent training, validation, and test sets, and employ accuracy as the performance metric. As described in Section 4.1, we generate a diverse set of 400 training prompts and 50 test prompts, ensuring that the test prompts are distinct from the training prompts, see the Appendix C for details. This separation rigorously evaluates the ability of model to generalize to unseen prompt formulations. We establish five baselines for comparison to isolate the impact of prompt engineering on fine-tuning: the pre-trained model without fine-tuning (Base Model); fine-tuning with human-designed prompts (User-Specified Prompt) as in Wei et al. (2024); fine-tuning with the prompt exhibiting the highest accuracy on the training set (Top-Accuracy Prompt); fine-tuning with the most robust prompt generated by BATprompt Shi et al. (2024) (BATprompt); and fine-tuning with the optimal prompt selected by ZOPO Hu et al. (2024) from the training prompt set (ZOPO Prompt). The key distinction between these methods lies in the prompt selection for fine-tuning. Critically, all models, including the baselines, are evaluated using the same set of 50 test prompts. This consistent evaluation protocol allows us to directly compare performance consistency and variation across methods. Our implementation leverages the Llama-factory framework (Zheng et al., 2024) and is evaluated using the Opencompass framework (Contributors, 2023). Detailed experimental configurations are provided in Appendix A. All experiments are conducted on NVIDIA A100, V100, 4090, and L40 GPUs to ensure efficient and scalable evaluation.
@@ 5.2 Main Results
PAFT demonstrates strong prompt robustness
As shown in Table 1, Figure 4, and Figure 6, PAFT exhibits remarkably low variance across all evaluation tasks, indicating excellent prompt robustness. Compared to other methods, PAFT achieves significantly lower variance, attributable to its unique dynamic prompt selection strategy. This strategy continuously adjusts the prompt during training, compelling the model to learn essential task features rather than overfitting to a specific prompt format. This contrasts sharply with the other baseline models. User-specified prompts rely on manually designed prompts, making it challenging to ensure both quality and diversity, especially without domain expertise. While TopAccuracy and ZOPO select the prompt exhibiting the highest accuracy on the training set, they are prone to overfitting to specific prompts and exhibit poor generalization. Although BATprompt also considers prompt robustness, its generated robust prompts are less effective than PAFT. In summary, the low variance of PAFT implies more stable performance and stronger generalization across diverse prompts, leading to higher reliability in practical applications. Specifically, models trained with PAFT can be used to develop more user-friendly question-answering systems, agent systems independent of input-output formats, and even to better decouple LLM capabilities from prompts, enabling more accurate LLM evaluation. PAFT achieves top performance on the majority of prompts, significantly outperforming all baselines (Table 1, Top column). Furthermore, PAFT maintains high training efficiency, A detailed discussion of training efficiency is provided in Appendix B.
@@
PAFT achieves state-of-the-art performance
As shown in Table 1, Figure 4, and Figure 6, PAFT achieves the highest average accuracy across all evaluated reasoning and reading comprehension tasks, significantly outperforming other baseline models. Specifically, PAFT surpasses other methods on tasks such as HellaSwag, PIQA, Winogrande, RACE, demonstrating its excellent performance across diverse natural language processing tasks. This superior performance stems from PAFT’s prompt robustness, enabling the model to better grasp the core essence of each task and maintain high performance across diverse prompt formulations. For instance, strong performance of PAFT on the open text generation task (HellaSwag) can be attributed to its dynamic prompt selection strategy, facilitating improved capture of contextual information. Its success on the physical common sense reasoning task (PIQA) can be attributed to its enhanced ability to utilize common sense knowledge. Similarly, its performance on the reference resolution task (Winogrande) can be attributed to its improved understanding of sentence structure and semantic relations, while its success on the reading comprehension task (RACE) can be attributed to its improved capture of topic and key information. In essence, this performance gain arises from PAFT’s decoupling of the prompt from the task itself, allowing the model to focus on learning the fundamental aspects of the downstream tasks.
PAFT enhances inference efficiency
In addition to robustness and performance, PAFT also significantly enhances inference efficiency. By fundamentally enhancing the ability of model to understand the core semantics of tasks, PAFT enables the model to solve problems more effectively, generating fewer tokens. This capability directly translates to faster inference speeds, as the model avoids redundant or unnecessary outputs and focuses on concise, accurate responses. To quantify this improvement, we measured the average end-to-end inference time across all test prompts and datasets, from the input prompt to the final output. As shown in Table 2, models trained with PAFT consistently achieve the fastest inference speeds compared to the baseline methods. This improvement is a direct result of PAFT’s inherent prompt robustness. By decoupling model performance from the specific prompt wording, PAFT operates consistently and efficiently regardless of the input prompt. In essence, PAFT promotes more effective generalization and eliminates the need for prompt-specific adaptation during inference. Additionally, our training regime covers a wide range of prompt wordings, avoiding the potential performance degradation or increased computation typically required to handle unexpected or unevenly distributed prompts during inference. This consistency and efficiency is especially valuable in real-world applications that require fast response times, such as dialogue systems or time-sensitive information retrieval. Our enhanced inference efficiency translates to a better user experience and reduced computational resources required for deployment, making it a more practical and scalable solution.
@@ 5.3 Ablation Studies
This ablation study demonstrates the robustness of PAFT to the hyperparameters (iterations per prompt) and (epochs). As shown in Table 3, PAFT achieves stable performance across a broad range of (1 to 8) and (3 to 6) values, with minimal fluctuations in accuracy and variance. Notably, PAFT achieves near-optimal performance with default settings (, ), attaining an average accuracy of 87.46%() across all tasks. This robustness reduces the need for extensive hyperparameter tuning, making PAFT a practical and efficient solution for real-world applications.
PAFT achieves strong performance with limited training prompts
@@
We conduct an ablation study to investigate the impact of varying numbers of training prompts on model performance, thus validating the effectiveness of PAFT. The experimental results, shown in Figure 5, demonstrate that as the number of prompts increases, the average accuracy of the model significantly improves, while the standard deviation decreases, indicating more stable and reliable performance. However, the performance gains diminish as the number of prompts increases, with only marginal improvements observed beyond a certain threshold. This suggests that while adding prompts can enhance performance, PAFT achieves competitive results with a minimal number of prompts, rendering excessive prompts unnecessary. In most cases, PAFT achieves strong performance with as few as 10 high-quality prompts, and further increases yield only marginal gains. The efficiency of PAFT is particularly notable, as it delivers excellent performance with a minimal number of prompts, making it highly suitable for resource-constrained scenarios where computational efficiency is critical. These findings underscore the practicality and efficiency of PAFT, offering a robust and efficient solution for real-world applications.
@@ 6 Conclusion
PAFT offers a compelling solution for enhancing the prompt robustness of LLMs. By dynamically adjusting prompts during fine-tuning, PAFT significantly improves model generalization and performance across diverse prompt formulations. Notably, PAFT boosts inference speed with maintained training cost. This approach paves the way for more reliable and efficient LLM deployment in real-world applications.
@@
Appendix A Experimental setting
@@ A.1 Dataset
A.2 Specific experimental parameters
@@
Appendix B Training cost and inference time
PAFT Maintains Training Efficiency
We now turn our attention to the training efficiency of PAFT. A critical consideration for any practical fine-tuning approach is its impact on training time. Introducing complex mechanisms or additional computational overhead can significantly hinder the training process, especially when dealing with large language models and extensive datasets. Therefore, it is essential to demonstrate that PAFT does not introduce such burdens.
@@ Appendix B Training cost and inference time
Efficient Candidate Prompt Generation
A key aspect of PAFT’s effectiveness lies in its ability to generate a diverse and high-quality set of candidate prompts efficiently. The process of constructing these candidate prompts involves leveraging the capabilities of external large language models (LLMs), which naturally raises the question of associated costs. Specifically, we sought to quantify the token usage required for candidate prompt generation, as this directly translates to the expense incurred when interacting with commercial LLM APIs.
@@
Appendix C Prompt
In this section, we present a selection of training and test prompts to illustrate the efficacy of our prompt construction algorithm and to provide a clearer understanding of operational process of PAFT. Due to space constraints, we only list 10 prompts as examples. Section C.1 showcases examples of training prompts, Section C.2 highlights test prompts, and Section C.3 outlines the prompts utilized by the baseline method.
C.1 Train prompt

### v2

@@ 5 Empirical Results
We evaluate our PAFT framework through comprehensive experiments. Sec. 5.1 describes datasets and experimental setup, Sec. 5.2 analyzes key findings, and Sec. 5.3 presents ablation studies examining critical framework components.
@@ 5.1 Datasets and Setup
Benchmark Selection. As a pioneering work addressing prompt robustness in LLMs through training, we conduct experiments across a diverse set of tasks and benchmarks to ensure a comprehensive evaluation. Our methods involve Supervised Fine-Tuning (SFT) and a reinforcement learning approach, GRPO. For our SFT experiments, we selected benchmarks to cover a wide range of capabilities. Specifically, we use HellaSwag Zellers et al. (2019) for knowledge understanding, WinoGrande Sakaguchi et al. (2019) for language understanding, and RACE Lai et al. (2017) for reading reasoning capabilities. For grounding and abstractive summarization, we employ PIQA Bisk et al. (2019). To evaluate more specialized skills, we utilize HumanEval Chen et al. (2021) for coding, T-Eval Chen et al. (2024) for tool use, and Xstory_cloze Lin et al. (2021) for multi-turn dialogues and multilingual tasks. For our GRPO experiments, we focus on mathematical reasoning. We use GSM8K Cobbe et al. (2021) for math reasoning capabilities and Geometry3k Lu et al. (2021) for multimodal mathematical reasoning.
Prompt Sets. As detailed in Section 4.1, we constructed distinct prompt sets for training and evaluation. The training set contains 400 diverse prompts generated exclusively via LLMs; this approach demonstrates that PAFT can fully automate the construction of training materials without manual intervention. For evaluation, we carefully designed a separate test set containing 50 prompts that includes not only LLM-generated prompts but also intentionally human-written instructions. The inclusion of human-written prompts is crucial for comprehensively validating the generalization of our model capabilities and its practical utility in real-world scenarios. This strict separation between a fully synthetic training set and a hybrid test set ensures a rigorous assessment of our method’s effectiveness. Further details on the prompt generation process are provided in Appendix D.
Baseline Comparisons. We establish five baselines to isolate the impact of prompt engineering on fine-tuning performance: the original pre-trained model (Base Model); the model fine-tuning with human-designed prompts (User) following Wei et al. (2024); the model fine-tuning with the highest accuracy of training prompts (TopAccuracy); the model fine-tuning with BATprompt Shi et al. (2024) most robust prompt (BATprompt); and fine-tuning with ZOPO Hu et al. (2024) optimal prompt selection (ZOPO). All models, including baselines, are evaluated on identical test prompts, enabling direct comparison of performance consistency across methods.
Experimental Setup. To comprehensively evaluate the effectiveness of PAFT, our experiments cover two main training paradigms: supervised fine-tuning (SFT) and reinforcement learning fine-tuning (RLFT). For the SFT paradigm, we adopt Low-Rank Adaptation (LoRA Hu et al. (2022)) as a representative method, and the specific experimental parameters are shown in the Appendix B.2; for RLFT, we employ Group Relative Policy Optimization (GRPO Shao et al. (2024)) and the specific experimental parameters are shown in the Appendix B.3. In these settings, we utilize a series of large language models (LLMs), including Llama3-8B, Llama-3.1-8B, Llama-3.2-3B Meta (2024), Qwen2.5-7B, and Qwen2.5-VL-7B Qwen et al. (2025). Detailed correspondence between model datasets and training paradigms is provided in Appendix Table 7. Our implementation is based on the Llama-factory framework, and all evaluations are performed using OpenCompass. All experiments are conducted on NVIDIA A100, V100, 4090, and L40 GPU clusters. For detailed configuration, see the Appendix B.
@@ 5.2 Main Results
Prompt Robustness. As demonstrated across Tables 1, Table 2, Figures 4, 5, 6, and 7, PAFT exhibits remarkably low variance across all evaluation tasks, indicating superior prompt robustness. This enhanced stability stems from our dynamic prompt selection strategy (Sec. 4.2), which continuously adjusts prompts during training, compelling the model to learn essential task features rather than overfitting to specific prompt formats. In contrast, baseline approaches face significant limitations: user prompts rely on manual design with inconsistent quality; TopAccuracy and ZOPO tend to overfit to high-performing training prompts with poor generalization; and while BATprompt addresses robustness, it remains less effective than our method. The low variance of PAFT translates to more stable performance and stronger generalization across diverse prompts, enabling development of more user-friendly QA systems, format-independent agent systems, and directly evaluate the true ability of LLMs by better decoupling the ability from the prompting engineering. Notably, PAFT achieves acceptable performance across most prompts, significantly outperforming all baselines (Table 1, Top column) while maintaining high training efficiency (detailed in Appendix C).
| SFT Model | PAFT Model | Improvement | ||||
| Dataset | Min | Con | Min | Con | Min | Con |
| HellaSwag | 87.20 | 61.26 | 91.30 | 84.61 | +4.10 | +23.35 |
| PIQA | 75.16 | 62.13 | 88.72 | 84.12 | +13.56 | +22.99 |
| HumanEval | 45.26 | 12.46 | 52.89 | 50.16 | +7.63 | +37.70 |
| RACE-mid | 72.36 | 50.16 | 85.07 | 83.27 | +12.71 | +33.11 |
| RACE-high | 71.68 | 49.67 | 84.26 | 81.26 | +12.58 | +31.59 |
| GSM8K | 40.36 | 10.26 | 75.13 | 74.13 | +34.77 | +63.87 |
| Geometry3k | 31.26 | 12.37 | 38.90 | 37.12 | +7.64 | +24.75 |
| T-Eval | 42.13 | 19.26 | 61.27 | 59.17 | +19.14 | +39.91 |
To quantify this robustness under more demanding conditions, we introduce two stringent metrics. First, we measure minimum accuracy (Min) on the test set of 50 unseen prompts to evaluate worst-case performance under normal conditions. Second, we assess conditional accuracy (Con) on a challenging set of 10 adversarially crafted prompts containing paraphrases, misspellings, or other modifications to measure resilience against noisy inputs. As shown in Table 4, the PAFT-trained model achieves significantly higher minimum and conditional accuracy across all datasets. This result underscores the ability of PAFT to maintain effective performance even when faced with substantial prompt perturbations and adversarial noise.
SOTA Performance. As demonstrated across our experiments (Tables 1 and 2; Figures 4–7), PAFT consistently achieves SOTA performance by significantly outperforming existing baselines. This superior performance stems directly from our Dynamic Fine-Tuning algorithm (Algorithm 1; Section 4.2), which effectively decouples the underlying fundamental principles of a task from any specific prompt formulation. This decoupling allows the model to focus on learning the essential features of task, rather than becoming entangled in the nuances of a particular prompt. Ultimately, this process enables the model to learn the fundamental principles of downstream tasks instead of merely overfitting to superficial prompt patterns, which is the key reason why PAFT model can achieve its SOTA generalization and performance.
@@
Inference Efficiency. PAFT enhances inference efficiency not by accelerating per-token generation speed, but by enabling the model to produce correct and concise responses using significantly fewer tokens. Our measurements across all test prompts and datasets (Table 3) demonstrate that this token reduction leads to consistently faster overall inference times compared to baseline methods. This efficiency stems from the prompt robustness of model, which is a core outcome of our training approach. Unlike baseline models that may overfit to superficial prompt patterns, PAFT learns underlying task principles. As a result, even in the face of changing or modified prompts, the model remains focused on the fundamental task objectives, avoiding the need to output redundant dialogue content, such as understanding unknown instructions. This making PAFT particularly valuable for real-world applications requiring rapid responses, such as dialogue systems and AI agents, while simultaneously reducing computational resource requirements. See Appendix C.2 for more detail.
@@ 5.3 Ablation Studies
Hyperparameter Robustness. This ablation study demonstrates the robustness of PAFT to the hyperparameters (iterations per prompt) and (epochs). As shown in Table 5, PAFT achieves stable performance across a broad range of (1 to 8) and (3 to 6) values, with minimal fluctuations in accuracy and variance. Notably, PAFT achieves near-optimal performance with default settings (, ), attaining an average accuracy of 87.46%() across all tasks. This robustness reduces the need for extensive hyperparameter tuning, making PAFT a practical and efficient solution for real-world applications.
@@
Impact of Training Prompt Quantity. We conduct an ablation study to investigate the impact of varying numbers of training prompts on model performance, thus validating the effectiveness of PAFT. The experimental results, shown in Figure 8, demonstrate that as the number of prompts increases, the average accuracy of the model significantly improves, while the standard deviation decreases, indicating more stable and reliable performance. However, the performance gains diminish as the number of prompts increases, with only marginal improvements observed beyond a certain threshold. This suggests that while adding prompts can enhance performance, PAFT achieves competitive results with a minimal number of prompts, rendering excessive prompts unnecessary. In most cases, PAFT achieves strong performance with as few as 10 high-quality prompts, and further increases yield only marginal gains. The efficiency of PAFT is particularly notable, as it delivers excellent performance with a minimal number of prompts, making it highly suitable for resource-constrained scenarios where computational efficiency is critical. These findings underscore the practicality and efficiency of PAFT, offering a robust and efficient solution for real-world applications.
6 Theoretical Insights
The capability of PAFT to generalize effectively to unseen prompt formulations can be rigorously understood through the lens of domain adaptation theory Ben-David et al. (2006, 2010). In this theoretical construct, the collection of training prompts along with the task-specific training data delineates the source domain. Besdies, the set of novel test prompts , paired with , represents the target domain. PAFT aims to learn a model , where denotes the hypothesis class, by minimizing the empirical risk computed over instances where each prompt is sampled from .
A foundational result from domain adaptation theory Ben-David et al. (2010) provides an upper bound on the expected risk of on the target prompt distribution with :
Here, is the empirical risk on training prompts. The term signifies model complexity (e.g., related to Rademacher complexity Yin et al. (2020)), which typically diminishes as the number of distinct training prompts increase; this term captures the generalization gap on the source domain. The divergence between the training and test prompt distributions is quantified by . Finally, encapsulates the optimal joint error achievable by a hypothesis in on both domains. The key of PAFT is designed to optimize this bound for improved generalization.
Complexity Control. By employing a substantial number of distinct training prompts , PAFT inherently works to reduce the complexity term . This ensures that the model performance observed on the training prompts becomes a more faithful estimator of its true performance across the entire distribution, fostering more stable learning. This effect is empirically supported by our ablation studies in Section 5.3 (Figure 8), which demonstrate improved stability with more prompts.
Domain Alignment. Minimizing the domain discrepancy term is critically dependent on constructing a diverse and comprehensive set of candidate prompts (see Section 4.1 for details). A more diverse is more likely to effectively cover or closely approximate the diverse and unseen distribution of test prompts . This approximation reduces the divergence between the training and test prompt distributions and enhances the transferability of knowledge learned from to .
Proposition 1 (MMD as an Upper Bound on Discrepancy).
The discrepancy term can be bounded by the Maximum Mean Discrepancy (MMD) Gao et al. (2021a) upper bound:
where MMD is defined as:
The proof of Proposition 1 is provided in the Appendix A. Therefore, we can quantify the upper bound of domain difference using MMD. As illustrated in Figure 9, an increasing number of diverse training prompts cover a wider semantic space, bringing closer to . This proximity reduces the upper bound of the target prompt distribution.
Generalization Guarantee. By minimizing the empirical risk across a sufficiently large and varied corpus of prompts, PAFT encourages the model to internalize the underlying task semantics, rather than merely memorizing superficial prompt structures. This principled approach is key to improving the model performance when confronted with novel and unencountered prompts.
@@ 7 Conclusion
PAFT offers a compelling solution for enhancing the prompt robustness of LLMs. By dynamically adjusting prompts during fine-tuning, PAFT significantly improves model prompt robustness and performance across diverse prompt formulations. Notably, PAFT boosts inference speed with maintained training cost. This approach paves the way for more reliable and efficient LLM deployment in real-world applications.
Acknowledgement
This work is supported in part by Shenzhen Science and Technology Program under Grant ZDSYS20220527171400002, the National Natural Science Foundation of China (NSFC) under Grants 62271324, 62231020 and 62371309.
@@ Acknowledgement
Appendix A Theoretical Proof
We first state a core assumption regarding the richness of the reproducing kernel Hilbert space (RKHS).
Assumption 1 (Richness of RKHS).
We assume that the RKHS generated by the chosen kernel function is large enough to contain all 1-Lipschitz functions. Formally, we assume that there exists a constant such that any function with a Lipschitz constant at most 1 (i.e., ) is contained in and has a bounded norm, i.e., .
Proof Sketch.
The difference term is the supremum over all 1-Lipschitz functions. According to the assumption 1, we can extend this function space to the RKHS sphere of radius , which contains all 1-Lipschitz functions, and then obtain the definition of MMD by rescaling.
This derivation formally shows that we can use MMD to quantize . ∎
Appendix B Experimental setting
@@ B.1 Dataset
| Type | Method | LLM | Dataset |
| Knowledge Understanding | SFT | Llama3-8B | HellaSwag |
| Language Understanding | SFT | Llama3-8B | WinoGrande |
| Math Reasoning Capabilities | GRPO | Qwen2.5-7B | GSM8K |
| Reading Reasoning Capabilities | SFT | Llama3-8B | RACE |
| Grounding and Abstractive Summarization | SFT | Llama3-8B | PIQA |
| Coding Capabilities | SFT | Qwen2.5-7B | HumanEval |
| Tool use | SFT | Llama-3.1-8B | T-Eval |
| Multi-turn dialogues and multilingual tasks | SFT | Llama-3.2-3B | Xstory_cloze |
| Multimodal mathematical reasoning | GRPO | Qwen2.5-VL-7B | Geometry3k |
B.2 Specific SFT experimental parameters
@@
B.3 PAFT Integration with GRPO
To demonstrate the versatility of our PAFT framework beyond SFT, we also integrated it with the Reinforcement Learning Fine-Tuning (RLFT) paradigm. We selected Group Relative Policy Optimization (GRPO) as a representative RLFT method and applied our dynamic prompting strategy to its training process for the GSM8K dataset. The setup is as follows:
Candidate Prompt Construction. Following the core principle of PAFT, we first constructed a diverse candidate prompt set for mathematical reasoning. We utilized multiple large language models (LLMs) to generate a wide variety of prompts related to math problems, which formed the prompt training set. To ensure a rigorous evaluation, additional prompts from real dialogues and other synthetic sources were used to create a distinct test set, guaranteeing that the prompt training and test sets were entirely distinct.
Dynamic GRPO Training Process. We integrated our dynamic sampling mechanism directly into the GRPO training loop. In each step of the standard GRPO process, instead of using a fixed instruction, we randomly sampled a prompt from our candidate training set. This sampled prompt (e.g., “Please help me solve this math problem {GSM8K_problem}”) was then combined with a problem instance from the GSM8K dataset to form the final input for the model’s computation step. This process aligns with the standard procedure for Soft Fine-Tuning.
Hyperparameters. The key hyperparameters used for our GRPO experiments are detailed in Table 9.
| Parameter | Value |
| Learning Rate | 5e-6 |
| Num Generations | 16 |
| Epochs | 10 |
Appendix C Training cost and inference time
C.1 Training cost
PAFT Maintains Training Efficiency. We now turn our attention to the training efficiency of PAFT. A critical consideration for any practical fine-tuning approach is its impact on training time. Introducing complex mechanisms or additional computational overhead can significantly hinder the training process, especially when dealing with large language models and extensive datasets. Therefore, it is essential to demonstrate that PAFT does not introduce such burdens.
@@ C.1 Training cost
Efficient Candidate Prompt Generation. A key aspect of PAFT’s effectiveness lies in its ability to generate a diverse and high-quality set of candidate prompts efficiently. The process of constructing these candidate prompts involves leveraging the capabilities of external large language models (LLMs), which naturally raises the question of associated costs. Specifically, we sought to quantify the token usage required for candidate prompt generation, as this directly translates to the expense incurred when interacting with commercial LLM APIs.
@@
C.2 Inference Times
In this section, we show some special cases to demonstrate that PAFT improves reasoning efficiency by improving the model’s understanding of the core task semantics, enabling concise and accurate responses with fewer tokens.
Appendix D Prompt
In this section, we present a selection of training and test prompts to illustrate the efficacy of our prompt construction algorithm and to provide a clearer understanding of operational process of PAFT. Due to space constraints, we only list 10 prompts as examples. Section D.1 shows how we guide LLMs to generate candidate prompts. Section D.2 showcases examples of training prompts, Section D.3 highlights test prompts, and Section D.4 outlines the prompts utilized by the baseline method.
D.1 Automated Prompt Generation Strategy
To construct a diverse and high-quality set of candidate prompts, we employ a strategy that leverages large language models (LLMs) through two distinct templating approaches: zero-shot and one-shot prompting. These templates are designed to be general, requiring only minor modifications to synthesize prompt sets for various datasets.
Zero-shot Prompting. Our zero-shot approach uses a general template that instructs an LLM to generate multiple prompt variations for a given task type, without being constrained by a specific problem instance. This method is effective for tasks where the output format is straightforward. For example, to generate prompts for a commonsense reasoning task like PIQA, we use the following instruction:
One-shot Prompting. For tasks requiring a specific output format, such as the step-by-step reasoning in mathematical problems, we use a one-shot template. This template provides the LLM with an explicit example of the desired output structure in addition to the generation instructions, thereby guiding the model’s response format[cite: 1]. For datasets like GSM8K and Geometry3K, our one-shot instruction is as follows:
This templating strategy is designed to produce prompts that are both adaptable and general. By crafting instructions that elicit the necessary information without being overly task-specific, we ensure the generated prompts can be applied across different datasets with minimal modification. This approach is fundamental to our goal of enhancing prompt robustness and practical applicability, demonstrating that our framework can automate the creation of effective and varied training prompts.
D.2 Train prompt